OpenStudy (anonymous):

Find the cost of 15 shares of stock at 13 1/2 per share.

OpenStudy (anonymous):

I'm still not sure about how you got that. DOd you just multiply 15 by 13?

OpenStudy (anonymous):

I meant do you just multiply 15 by 13?

OpenStudy (anonymous):

It is 13 and a half. So, it is 13.5

OpenStudy (anonymous):

Oh!! So you multiplied 15 times 13.5

OpenStudy (diyadiya):

\[13 \frac{1}{2}= 13+ \frac{1}{2}= \frac{13 \times 2 +1}{2}= \frac{27}{2}\]

OpenStudy (anonymous):

I'm sorry can you take it step by step? I still am not understanding a part yet

OpenStudy (anonymous):

Sorry...I messed up. I meant to do 27/2. Oooooooppppppssiiiiieeeee

OpenStudy (anonymous):

So... what..? :)

OpenStudy (diyadiya):

\[1share = 13 \frac{1}{2}= 27/2 \]\[15 shares = 15 \times 27/2\]

OpenStudy (diyadiya):

You want help with converting improper fraction into proper fraction?

OpenStudy (anonymous):

I have the answer and it's $202.50 but I'm not sure how they got this...

OpenStudy (diyadiya):

Did you understand how 13 1/2 became 27/2?

OpenStudy (anonymous):

no. :)

OpenStudy (diyadiya):

\[13 \frac{1}{2} \] is an improper fraction so you have to convert this before doing the problem

OpenStudy (diyadiya):

\[13 \frac{1}{2}= 13+ \frac{1}{2}\] ok?

OpenStudy (anonymous):

so thats 13.5. right:)

OpenStudy (anonymous):

Simpler to think in decimals may be? 13 and half = 13.5.

OpenStudy (diyadiya):

Yes you can simplify it to decimal as @GT said :)

OpenStudy (anonymous):

So then what? :)

OpenStudy (anonymous):

Multiply 13.5 by 15. You get 202.5.

OpenStudy (anonymous):

I get it now! Thank you! :D

OpenStudy (diyadiya):

Its given for 1 share its 13.5 so for 15 share 15 *13.5 =202.5

OpenStudy (anonymous):